# BF-STVSR: B-Splines and Fourier—Best Friends for High Fidelity Spatial-Temporal Video Super-Resolution

### v1

@@ Abstract
Enhancing low-resolution, low-frame-rate videos to high-resolution, high-frame-rate quality is essential for a seamless user experience, motivating advancements in Continuous Spatial-Temporal Video Super Resolution (C-STVSR). While prior methods employ Implicit Neural Representation (INR) for continuous encoding, they often struggle to capture the complexity of video data, relying on simple coordinate concatenation and pre-trained optical flow network for motion representation. Interestingly, we find that adding position encoding, contrary to common observations, does not improve—and even degrade—performance. This issue becomes particularly pronounced when combined with pre-trained optical flow networks, which can limit the model’s flexibility. To address these issues, we propose BF-STVSR, a C-STVSR framework with two key modules tailored to better represent spatial and temporal characteristics of video: 1) B-spline Mapper for smooth temporal interpolation, and 2) Fourier Mapper for capturing dominant spatial frequencies. Our approach achieves state-of-the-art PSNR and SSIM performance, showing enhanced spatial details and natural temporal consistency. Our code will be available soon.
@@ 1 Introduction
Enhancing low-resolution, low-frame-rate videos to high-resolution, high-frame-rate quality is crucial for delivering seamless user experiences. To address this, deep learning approaches for Video Super-Resolution (VSR) [3, 4, 43, 46, 44] and Video Frame Interpolation (VFI) [1, 10, 9, 13, 18, 38, 8, 50, 47] have been extensively studied. VSR typically enhances spatial resolution of target frames by leveraging information from neighboring frames, while VFI improves temporal resolution by predicting inherent motion in video data. However, many existing methods are limited by fixed scaling factors set during training, which restricts their flexibility in real-world applications.
@@
Figure 1: Illustration of BF-STVSR and results. (a) BF-STVSR captures the high-frequency spatial feature by Fourier Mapper and interpolates temporal information smoothly via B-spline Mapper. (b) Temporal visualization of the interpolated video. We visualize the changes of the interpolated frames over time $t$ for a selected x-axis (indicated by yellow vertical line in (a)).
@@
On the other hand, Implicit Neural Representation (INR) has recently garnered attention for its capability to represent signals continuously through a multi-layer perceptron (MLP), making it a promising approach for super-resolution (SR) tasks [5, 22, 32, 15]. Building on these advancements, recent studies have extended INR to video data to achieve Continuous Spatial-Temporal Video Super Resolution (C-STVSR), which enables spatial and temporal interpolation simultaneously at arbitrary scales [7, 6]. VideoINR [7] was the first method to map spatiotemporal coordinates ${(x,y,t)}$ to backward motion field, facilitating backward warping of spatial features to any temporal coordinate. MoTIF [6] improved on this by replacing the backward warping with forward warping, using softmax splatting [31]. In addition, to facilitate the learning in an explicit way, MoTIF supply forward optical flow maps estimated between reference frames as contextual information, using RAFT [42].
@@
While VideoINR and MoTIF successfully integrate INR into the C-STVSR task, they have notable limitations. Specifically, they generate target features by encoding latent features that are simply concatenated with target coordinates, without employing advanced position encoding techniques. This is surprising, given that various position encoding methods—such as Fourier encoding [27, 41]—are well-established and widely used in tasks like image SR with INR due to their effectiveness, having become a conventional process [22, 16, 32, 48]. This straightforward approach may fall short in capturing the nuanced details of spatial and temporal features, especially for motion features, which are inherently complex and dynamic. Consequently, both models struggle to retain high-frequency information in the encoded spatial features, a well-known limitation referred to as spectral bias [34, 41], resulting in the generation of lower-quality frames.
@@
Interestingly, however, we find that simply adding position encoding does not improve—and even degrade—performance in these models, an unexpected outcome that contrasts with the general success of position encoding in enhancing INR applications [45, 12, 28, 19]. This issue becomes particularly pronounced when combined with pre-trained optical flow networks. We conjecture that, while these networks provide useful guidance for motion representation, integrating them with position encoding can inadvertently limit the model’s flexibility to fully leverage diverse video information.
@@
To address these limitations, we propose BF-STVSR, a framework consisting of two modules: B-spline Mapper (BM) and Fourier Mapper (FM), each designed to handle temporal and spatial features. First, B-spline Mapper (BM) utilizes B-spline function, well-known established method for constructing smooth curves or surfaces [32]. This approach well-suited for capturing the continuous nature of video motion. Next, Fourier Mapper (FM) represents spatial features by estimating dominant frequency information of input video frames, effectively capturing fine details. Additionally, based on our earlier analysis, rather than directly integrating a pre-trained optical flow network into our framework, we incorporate it implicitly as a loss function during training to guide the network. This design allows the model to fully leverage the rich information inherent in video data, yielding enhanced interpolated outputs.
@@
In summary, our contributions are as follows: (1) We propose BF-STVSR, a framework that independently address the spatial and temporal axes by leveraging their distinct characteristics. To achieve this, (2) we design two dedicated components, B-spline Mapper (BM) for temporal motion representation and Fourier Mapper (FM) for spatial feature representation. (3) We achieve state-of-the-art performance in C-STVSR, demonstrating the effectiveness of our approach through extensive experiments.
@@
Figure 2: Schematic overview of our BF-STVSR. (a) First, two input frames are encoded as low-resolution feature maps. Based on these features, Fourier Mapper predicts the dominant frequency information, while B-spline Mapper predicts smoothly interpolated motion, which is then processed into optical flows at an arbitrary time $t$. The frequency information is temporally propagated by being warped with the optical flows. Finally, the warped frequency information is decoded to generate high-resolution interpolated RGB frame. (b) Fourier Mapper estimates the dominant frequency and its amplitude to capture fine-detail information from the given frames. (c) B-spline Mapper estimates B-spline coefficients to model inherent motion, which smoothly interpolates motion features temporally.
@@ 2 Related Work
2.1 Arbitrary Single Image Super Resolution
Single Image Super Resolution (SISR) methods [23, 24, 37, 54] have achieved impressive performance, but their reliance on fixed scales limits their applicability in real-world scenarios. To address this, several studies have proposed methods to perform super resolution at arbitrary scales [5, 51, 26, 22, 32]. LIIF [5] introduced an Implicit Neural Representation (INR) approach for super resolution, representing images continuously through local implicit functions and enabling arbitrary scale upsampling. [51, 26] further used position encoding to address the spectral bias problem [34]. Recently, LTE [22] proposed identifying dominant Fourier bases from latent features to effectively capture fine details and address spectral bias. Similarly, BTC [32] employed B-spline bases instead of Fourier bases to mitigate the Gibbs phenomenon observed in Screen Content Image Super Resolution. Inspired by these methods, we explore effective position encoding techniques for C-STVSR task, which reflect the characteristics of video data.
2.2 Spatial-Temporal Video Super Resolution
While conventional Video Super-Resolution (VSR) [2, 36, 3, 4] and Video Frame Interpolation (VFI) [10, 13, 18, 8, 9, 21, 38, 30, 31, 11, 20, 33, 35, 53] perform interpolation along either spatial or temporal axis, Spatial-Temporal Video Super Resolution (STVSR) conducts interpolation along both axes. Haris et al. [14] have introduced a unified framework for addressing STVSR and Xiang et al. [47] have proposed to use bidirectional deformable ConvLSTM. Although these studies demonstrate impressive performance in STVSR, they both have the limitation of only addressing STVSR at fixed scales. Recently, two works [7, 6] have been proposed for Continuous Spatial-Temporal Video Super Resolution (C-STVSR), which performs interpolation at arbitrary scales along both spatial and temporal axes. VideoINR [7] is the first work on C-STVSR, which takes space-time coordinates as input and maps the corresponding RGB value in continuous manner using INR. Following this, MoTIF [6] generates temporal features using optical flow and performs forward warping to predict the interpolated high-resolution frame features. Although these studies effectively tackle the C-STVSR task, relying solely on MLPs for spatial and temporal modeling leads to difficulties in learning the characteristics of the video. In this work, we adopt Fourier basis and B-spline basis functions to model spatial and temporal features of video data to address the aforementioned difficulties.
@@ 3.1 Overview
The overall flow of our method, BF-STVSR, is illustrated in Figure 2 (a). Our framework is built upon the pipeline of MoTIF [6]. Given two low-resolution frames $I_{0}^{L},I_{1}^{L}\in\mathbb{R}^{3\times H\times W}$, our goal is to generate a high-resolution intermediate frame $I_{t}^{H}\in\mathbb{R}^{3\times sH\times sW}$ at any time $t\in[0,1]$ with an arbitrary scale $s$. The encoder $E$ first takes the low-resolution frames as input and produces three latent features: $F_{0}^{L},F_{(0,1)}^{L},F_{1}^{L}\in\mathbb{R}^{C\times H\times W}$. Here, $F_{0}^{L}$ and $F_{1}^{L}$ represent the latent features of $I_{0}^{L}$ and $I_{1}^{L}$, while $F_{(0,1)}^{L}$ serves as a template feature for the intermediate frame, incorporating information from both input frames. Next, the latent features $F_{0}^{L},F_{1}^{L}$ are processed by two mappers. The B-spline Mapper (Section 3.2) predicts high-resolution optical flows to target time $t$, producing $M_{0\rightarrow t}^{H},M_{1\rightarrow t}^{H}\in\mathbb{R}^{2\times sH\times sW}$, while the Fourier Mapper (Section 3.3) estimates high-resolution spatial features at target scale $s$, resulting in $F_{0}^{H},F_{1}^{H}\in\mathbb{R}^{C\times sH\times sW}$. Finally, the high-resolution features $F_{0}^{H},F_{1}^{H}$ are temporally propagated to the target time $t$ using forward warping based on the predicted optical flow $M_{0\rightarrow t}^{H},M_{1\rightarrow t}^{H}$, generating intermediate features $F_{t}^{H}$. These warped features are then concatenated with target time $t$ and $F_{(0,1)}^{H}$, a nearest-neighbor upsampled version of $F_{(0,1)}^{L}$, and decoded to produce the high-resolution intermediate frame $I_{t}^{H}$.
@@ 3.2 Temporal B-spline Mapper
Accurately predicting the motion inherent in a video is key to generating a visually plausible intermediate frame from two neighboring frames. While various techniques have been proposed in video frame interpolation (VFI) studies [18, 8, 31, 11, 20, 33, 35, 53] to predict motion accurately, most of them are designed for fixed scale interpolation. Therefore, applying these methods to the C-STVSR task, which requires handling arbitrary scales, is not straightforward. To overcome this challenge, VideoINR [7] and MoTIF [6] use implicit neural representations (INR) with MLPs that take spatiotemporal coordinates as input, enabling motion modeling at arbitrary target times $t$ and scales $s$. While INR-based motion modeling provides flexibility in motion prediction, we find that it often struggles to capture the complex, dynamic nature of motion in video.
To better represent inherent motion, we introduce B-spline Mapper, which leverages the B-spline representation. B-spline bases are widely known for their effectiveness in modeling continuous signals [32], making them well-suited for capturing smooth, continuous motion in videos, where objects move smoothly and continuously, rather than in jerky manner. The detailed process of B-spline Mapper is described in Figure 2 (b). We modify the Space-Time Local Implicit Neural Functions (ST-INF) from MoTIF [6], resulting in our B-spline Mapper. Similar to ST-INF, B-spline Mapper predicts high-resolution forward motion vectors $M_{0\rightarrow t}^{H},M_{1\rightarrow t}^{H}$ and reliability maps $Z_{0\rightarrow t}^{H},Z_{1\rightarrow t}^{H}$ at arbitrary time $t\in[0,1]$. A key difference is that our B-spline Mapper takes encoded features $F_{0}^{L},F_{1}^{L}$ as input, rather than optical flows from an external network (e.g., RAFT [42]).
In addition, rather than directly predicting motion vectors to the target time $t$, our B-spline Mapper $p_{\psi}$ models the video’s inherent motion by predicting the B-spline coefficients and knots, as described in the following equation:
@@
Here, $c_{r}=p_{c}(z_{r},\delta_{r})$, $k_{r}=p_{k}(z_{r},\delta_{r})$, and $d=p_{d}(g)$. Specifically, $z_{r}=F_{t_{r}}^{L}(q_{r})$ is the latent feature vector at the point $q_{r}=(x_{r},y_{r})$, nearest to the query coordinates $q=(x,y)$, with the reference frame time index $t_{r}\in\{0,1\}$. The functions $p_{c}$, $p_{k}$, and $p_{d}$ are the estimators for the coefficients ($\mathbb{R}^{C+2}\mapsto\mathbb{R}^{C}$), knots ($\mathbb{R}^{C+2}\mapsto\mathbb{R}^{C}$), and dilation ($\mathbb{R}^{1}\mapsto\mathbb{R}^{C}$), respectively. $\hat{t}=|t-t_{r}|$ represents the relative temporal distance of the predicted feature to the reference frame, and $\delta_{r}(=q-q_{r})$ is the spatial relative coordinate between the query and reference coordinates. Finally, $g$ is the frame interval of the input video.
After linearly projecting the predicted B-spline representation using $f_{\theta_{b}}$, we obtain the motion vector $Z_{t_{r}\rightarrow t}^{H}(q)$ and reliability map $M_{t_{r}\rightarrow t}^{H}(q)$ at the query coordinates $q$:
@@
Using the predicted motion vectors, the spatial features $F^{H}_{0}$, $F^{H}_{1}$ and reliability maps are propagated to the target time $t$ via forward warping using softmax splatting [31]. Finally, we obtain intermediate latent feature $F_{t}^{H}$ and corresponding reliability map $Z_{t}^{H}$. By directly learning the underlying motion from the input frames instead of individually predicting each arbitrary time $t$, our B-spline Mapper provides a more robust and flexible motion modeling approach. Additionally, since our method does not rely on optical flow inference during the inference stage, it is more efficient compared to previous approaches like MoTIF [6].
Table 1: Performance comparison on the Fixed-scale STVSR baselines on Vid4, Gopro, and Adobe240 datasets. Results are evaluated using PSNR (dB) and SSIM metrics. All frames are interpolated by a factor of ×$\times 4$ in the spatial axis and ×$\times 8$ in the temporal axis. “Average” refers to metrics calculated across all 8 interpolated frames, while “Center” refers to metrics measured using $1^{st}$, $4^{th}$ and $9^{th}$ (that is single-frame interpolation) frames of the interpolated sequence. Red and blue indicate the best and the second best performance, respectively.
@@
Vid4 Gorpo-Center Gopro-Average Adobe-Center Adobe-Average
@@
Ours 25.80 / 0.7754 31.14 / 0.8893 30.20 / 0.8799 30.84 / 0.8877 30.14 / 0.8808 12.68
Table 2: Performance comparison on the C-STVSR baselines for out-of-distribution scale on Gopro dataset. Results are evaluated using PSNR (dB) and SSIM metrics. All frames are interpolated by a scaling factor specified on the table and metrics calculated across all interpolated frames. Bold indicates the best performance.
Temporal Scale
Spatial Scale
SuperSloMo [17] + LIIF [5] DAIN [1] + LIIF [5] TMNet [49] VideoINR [7] MoTIF [6] Ours
×$\times 8$$\times 4$ 28.83 / 0.8514 29.41 / 0.8669 30.04 / 0.8773 30.20 / 0.8799
×$\times 6$$\times 4$ 26.70 / 0.7988 26.71 / 0.7998 30.49 / 0.8861 30.78 / 0.8954 31.56 / 0.9064 31.68 / 0.9082
×$\times 6$$\times 6$ 23.47 / 0.6931 23.36 / 0.6902 - 25.56 / 0.7671 29.36 / 0.8505 29.44 / 0.8516
×$\times 6$$\times 12$ 21.92 / 0.6495 22.01 / 0.6499 - 24.02 / 0.6900 25.81 / 0.7330 25.78 / 0.7284
×$\times 12$$\times 4$ 25.07 / 0.7491 25.14 / 0.7497 26.38 / 0.7931 27.32 / 0.8141 27.77 / 0.8230 28.06 / 0.8287
×$\times 12$$\times 6$ 22.91 / 0.6783 22.92 / 0.6785 - 24.68 / 0.7358 26.78 / 0.7908 27.06 / 0.7961
×$\times 12$$\times 12$ 21.61 / 0.6457 21.78 / 0.6473 - 23.70 / 0.6830 24.72 / 0.7108 24.87 / 0.7096
×$\times 16$$\times 4$ 24.42 / 0.7296 24.20 / 0.7244 24.72 / 0.7526 25.81 / 0.7739 25.98 / 0.7758 26.40 / 0.7844
×$\times 16$$\times 6$ 23.28 / 0.6883 22.80 / 0.6722 - 23.86 / 0.7123 25.34 / 0.7527 25.81 / 0.7621
×$\times 16$$\times 12$ 21.80 / 0.6481 22.22 / 0.6420 - 22.88 / 0.6659 23.88 / 0.6923 24.22 / 0.6950

### v2

@@ Abstract
While prior methods in Continuous Spatial-Temporal Video Super-Resolution (C-STVSR) employ Implicit Neural Representation (INR) for continuous encoding, they often struggle to capture the complexity of video data, relying on simple coordinate concatenation and pre-trained optical flow networks for motion representation. Interestingly, we find that adding position encoding, contrary to common observations, does not improve---and even degrades---performance. This issue becomes particularly pronounced when combined with pre-trained optical flow networks, which can limit the model’s flexibility. To address these issues, we propose BF-STVSR, a C-STVSR framework with two key modules tailored to better represent spatial and temporal characteristics of video: 1) B-spline Mapper for smooth temporal interpolation, and 2) Fourier Mapper for capturing dominant spatial frequencies. Our approach achieves state-of-the-art in various metrics, including PSNR and SSIM, showing enhanced spatial details and natural temporal consistency. Our code is available here.
@@ 1 Introduction
Enhancing low-resolution, low-frame-rate videos to high-resolution, high-frame-rate quality is crucial for delivering seamless user experiences. To address this, deep learning approaches for Video Super-Resolution (VSR) [2, 32, 3, 4] and Video Frame Interpolation (VFI) [13, 26, 28, 31, 45] have been extensively studied. VSR typically enhances spatial resolution of target frames by leveraging information from neighboring frames, while VFI improves temporal resolution by predicting inherent motion in video data. However, many existing methods are limited by fixed scaling factors determined during training, which restricts their adaptability to real-world applications.
@@
Figure 1: Illustration of BF-STVSR and results. (a) BF-STVSR captures the high-frequency spatial features by Fourier Mapper and interpolates temporal information smoothly via B-spline Mapper. (b) We visualize the changes of the interpolated frames over time $t$ for a selected x-axis (yellow vertical line in (a)).
@@
On the other hand, Implicit Neural Representation (INR) has recently garnered attention for its capability to represent signals continuously through a multi-layer perceptron (MLP), making it a promising approach for super-resolution (SR) tasks [5, 16, 27, 12]. Building on these advancements, recent studies have extended INR to video data to achieve Continuous Spatial-Temporal Video Super-Resolution (C-STVSR), which enables spatial and temporal interpolation simultaneously at arbitrary scales [7, 6]. VideoINR [7] was the first method to map spatiotemporal coordinates ${(x,y,t)}$ to backward motion field, facilitating backward warping of spatial features to any temporal coordinate. MoTIF [6] improved on this by replacing the backward warping with forward warping, using softmax splatting [26]. In addition, to facilitate the learning in an explicit way, MoTIF supply optical flow maps estimated between reference frames as contextual information, using the pre-trained optical flow network, RAFT [37].
@@
While VideoINR and MoTIF successfully integrate INR into the C-STVSR task, they have notable limitations. Specifically, they generate target features by encoding latent features that are simply concatenated with target coordinates, without employing advanced position encoding techniques. This simple coordinate concatenation may fall short in capturing the nuanced details of spatial and temporal features, especially for motion features, which are inherently complex and dynamic. Consequently, both models struggle to retain high-frequency information in the encoded spatial features, a well-known limitation referred to as spectral bias [30, 36], resulting in the generation of lower-quality frames. This is surprising, given that various position encoding methods---such as Fourier encoding [23, 36]---are well-established and widely used in tasks like image SR with INR due to their effectiveness, having become a conventional process [16, 17, 27, 41].
@@
Interestingly, however, we find that simply adding position encoding does not improve---and even degrades---performance in these models, an unexpected outcome that contrasts with the general success of position encoding in enhancing INR applications [38, 10, 24, 15]. This issue becomes particularly pronounced when combined with pre-trained optical flow networks. We conjecture that, while these networks provide useful guidance for motion representation, integrating them with position encoding can inadvertently limit the model’s flexibility to fully leverage diverse video information.
@@
To address these limitations, we propose BF-STVSR, a framework consisting of two modules: B-spline Mapper and Fourier Mapper, each designed to handle temporal and spatial features. First, B-spline Mapper utilizes B-spline basis functions, well-known established method for constructing smooth curves or surfaces [27]. This approach is well-suited for capturing the continuous nature of video motion. Next, Fourier Mapper represents spatial features by estimating dominant frequency information of input video frames, effectively capturing fine details. Additionally, unlike MoTIF [6], B-spline Mapper models motion directly from encoded video features instead of relying on a pre-trained optical flow network. This not only allows the encoder to retain richer motion information for more accurate motion estimation but also improves efficiency by eliminating the need for the additional optical flow computation. Furthermore, our approach maintains reliable performance even without incorporating a pre-trained optical flow guidance in the training objective, further simplifying the overall framework.
@@
In summary, our contributions are as follows: (1) We propose BF-STVSR, a framework consisting of two dedicated components, B-spline Mapper for temporal motion representation and Fourier Mapper for spatial feature representation, addressing the spatial and temporal axes independently. (2) BF-STVSR estimates motion directly from encoded video features, enhancing efficiency and simplifying the framework. (3) Our BF-STVSR achieves state-of-the-art performance on C-STVSR, demonstrating the effectiveness of our approach through extensive experiments.
@@
Figure 2: Schematic overview of our BF-STVSR. (a) First, two input frames are encoded as low-resolution feature maps. Based on these features, Fourier Mapper predicts the dominant frequency information, while B-spline Mapper predicts smoothly interpolated motion representation, which is then processed into motion vectors at an arbitrary time $t$. The frequency information is temporally propagated by being warped with the predicted motion vectors. Finally, the warped feature is decoded to generate high-resolution interpolated RGB frame. (b) Fourier Mapper estimates the dominant frequencies and their amplitude to capture fine-detail information from the given frames. (c) B-spline Mapper estimates B-spline coefficients to model inherent motion, which smoothly interpolates motion features temporally.
@@ 2 Related Work
2.1 Arbitrary Single Image Super-Resolution
Single Image Super-Resolution (SISR) methods [19, 20, 46] have achieved impressive performance, but their reliance on fixed scales limits their applicability in real-world scenarios. To address this, several studies have proposed methods to perform super-resolution at arbitrary scales [5, 22, 16, 27]. LIIF [5] introduced an Implicit Neural Representation (INR) for arbitrary scale image super-resolution, representing images continuously through local implicit functions. IPE [22] further used position encoding to address the spectral bias [30]. Recently, LTE [16] proposed identifying dominant Fourier bases from latent features to effectively capture fine details and address spectral bias. Similarly, BTC [27] employed B-spline bases instead of Fourier bases to mitigate the Gibbs phenomenon observed in Screen Content Image Super-Resolution. Inspired by these methods, we explore effective position encoding techniques for C-STVSR, which reflect the characteristics of video data.
2.2 Spatial-Temporal Video Super-Resolution
While conventional Video Super-Resolution (VSR) [2, 32, 3, 4] and Video Frame Interpolation (VFI) [13, 26, 28, 31, 45] perform interpolation along either spatial or temporal axis, Spatial-Temporal Video Super-Resolution (STVSR) conducts interpolation along both axes. Haris et al. [11] have introduced a unified framework for addressing STVSR and Xiang et al. [40] have proposed to use bidirectional deformable ConvLSTM. Although these studies demonstrate impressive performance in STVSR, they both have the limitation of only addressing STVSR at fixed scales. Recently, two works [7, 6] have been proposed for Continuous Spatial-Temporal Video Super-Resolution (C-STVSR), which enables interpolation at arbitrary scales along both spatial and temporal axes. VideoINR [7] is the first work on C-STVSR, which takes spatiotemporal coordinates as input and maps the corresponding RGB value in continuous manner using INR. Following this, MoTIF [6] generates temporal features using optical flows and performs forward warping to predict the interpolated high-resolution frame features. Although these studies effectively tackle the C-STVSR, relying solely on MLPs for spatial and temporal modeling leads to difficulties in learning the characteristics of the video. In this work, we adopt Fourier and B-spline basis functions to model spatial and temporal features of video data to address the aforementioned difficulties.
@@ 3.1 Overview
The overall flow of our method, BF-STVSR, is illustrated in Fig 2 (a). Our framework is built upon the pipeline of MoTIF [6], but differs in that it removes the need for an external optical flow network (e.g., RAFT [37]) by introducing a learnable internal motion modeling approach based on B-spline and Fourier Mappers. Specifically, given two low-resolution frames $I_{0}^{L},I_{1}^{L}\in\mathbb{R}^{3\times H\times W}$, our goal is to generate a high-resolution intermediate frame $I_{t}^{H}\in\mathbb{R}^{3\times sH\times sW}$ at any time $t\in[0,1]$ with an arbitrary scale $s$. The encoder $E$ first takes the low-resolution frames as input and produces three latent features: $F_{0}^{L},F_{(0,1)}^{L},F_{1}^{L}\in\mathbb{R}^{C\times H\times W}$. Here, $F_{0}^{L}$ and $F_{1}^{L}$ represent the latent features of $I_{0}^{L}$ and $I_{1}^{L}$, while $F_{(0,1)}^{L}$ serves as a template feature for the intermediate frame, incorporating information from both input frames. The latent features $F_{0}^{L}$, $F_{1}^{L}$ are processed by (1) the B-spline Mapper (Sec 3.2), which predicts high-resolution motion vectors $M_{0\rightarrow t}^{H},M_{1\rightarrow t}^{H}\in\mathbb{R}^{2\times sH\times sW}$ to the target time $t$, and (2) the Fourier Mapper (Sec 3.3), which estimates high-resolution spatial features $F_{0}^{H}$, $F_{1}^{H}\in\mathbb{R}^{C\times sH\times sW}$ at scale $s$. Finally, the high-resolution features $F_{0}^{H},F_{1}^{H}$ are temporally propagated to the target time $t$ using forward warping based on the predicted motion vectors $M_{0\rightarrow t}^{H},M_{1\rightarrow t}^{H}$, generating intermediate features $F_{t}^{H}$. These warped features are then concatenated with target time $t$ and $F_{(0,1)}^{H}$, a nearest-neighbor upsampled $F_{(0,1)}^{L}$, and decoded to produce the high-resolution intermediate frame $I_{t}^{H}$.
@@ 3.2 Temporal B-spline Mapper
Previous C-STVSR approaches [7, 6] employ implicit neural representations (INR) using MLPs that take spatiotemporal coordinates as input, enabling motion modeling at arbitrary target times $t$ and scales $s$. While INR-based motion modeling offers flexibility in motion prediction, we observe that it often struggles to effectively capture the complex and dynamic nature of motion in videos.
To better represent inherent motion, we introduce B-spline Mapper, which leverages the B-spline representation. B-spline bases are widely known for their effectiveness in modeling continuous signals [27], making them well-suited for capturing smooth, continuous motion in videos, where objects move smoothly and continuously, rather than in jerky manner. The detailed process of B-spline Mapper is described in Fig 2 (b). We modify the Space-Time Local Implicit Neural Functions (ST-INF) from MoTIF [6], resulting in our B-spline Mapper. Similar to ST-INF, B-spline Mapper predicts high-resolution forward motion vectors $M_{0\rightarrow t}^{H},M_{1\rightarrow t}^{H}$ and reliability maps $Z_{0\rightarrow t}^{H},Z_{1\rightarrow t}^{H}$ at arbitrary time $t\in[0,1]$. A key difference is that our B-spline Mapper takes encoded features $F_{0}^{L},F_{1}^{L}$ as input, rather than optical flows from an external network (e.g., RAFT [37]).
In addition, rather than directly predicting motion vectors to the target time $t$, our B-spline Mapper $p_{\psi}$ models the inherent motion in the video by predicting B-spline coefficients and knots, as described in the following equation:
@@
Here, $c_{r}=p_{c}(z_{r},\delta_{r})$, $k_{r}=p_{k}(z_{r},\delta_{r})$, and $d=p_{d}(g)$. Specifically, $z_{r}=F_{t_{r}}^{L}(q_{r})$ is the latent feature vector at the coordinate $q_{r}=(x_{r},y_{r})$, nearest to the query coordinates $q=(x,y)$, with the reference frame time index $t_{r}\in\{0,1\}$. The functions $p_{c}$, $p_{k}$, and $p_{d}$ are the estimators for the coefficients ($\mathbb{R}^{C+2}\mapsto\mathbb{R}^{C}$), knots ($\mathbb{R}^{C+2}\mapsto\mathbb{R}^{C}$), and dilation ($\mathbb{R}^{1}\mapsto\mathbb{R}^{C}$), respectively. $\hat{t}=|t-t_{r}|$ represents the relative temporal distance of the predicted feature to the reference frame, and $\delta_{r}(=q-q_{r})$ is the spatial relative coordinate between the query and reference coordinates. Finally, $g$ is the frame interval of the input video.
After linearly projecting the predicted B-spline representation using $f_{\theta_{b}}$, we obtain the motion vector $M_{t_{r}\rightarrow t}^{H}(q)$ and reliability map $Z_{t_{r}\rightarrow t}^{H}(q)$ at the query coordinates $q$:
@@
Using the predicted motion vectors, the spatial features $F^{H}_{0}$, $F^{H}_{1}$ and reliability maps are propagated to the target time $t$ via forward warping using softmax splatting [26]. Finally, we obtain intermediate latent feature $F_{t}^{H}$ and corresponding reliability map $Z_{t}^{H}$. By directly learning the underlying motion from the input frames instead of individually predicting each arbitrary time $t$, our B-spline Mapper provides a more robust and flexible motion modeling approach. Note that, since our method does not rely on an external optical flow network, it offers more efficient and self-contained solution compared to prior approaches like MoTIF [6].
Table 1: Performance comparison on the Fixed-scale STVSR baselines on Vid4, GoPro, and Adobe240 datasets. $\mathcal{L}_{RAFT}$ refers the optical flow supervision. Results are evaluated using PSNR (dB) and SSIM metrics. All frames are interpolated by a factor of ×$\times 4$ in the spatial axis and ×$\times 8$ in the temporal axis. ‘‘Average" refers to metrics calculated across all 8 interpolated frames, while ‘‘Center" refers to metrics measured using $1^{st}$, $4^{th}$ and $9^{th}$ (that is the center-frame interpolation) frames of the interpolated sequence. Red and blue indicate the best and the second best performance, respectively.
@@
Vid4 GoPro-Center GoPro-Average Adobe-Center Adobe-Average
@@
BF-STVSR + $\mathcal{L}_{RAFT}$ (Ours) 25.80 / 0.7754 31.14 / 0.8893 30.20 / 0.8799 30.84 / 0.8877 30.14 / 0.8808 13.47
BF-STVSR (Ours) 25.85 / 0.7772 31.17 / 0.8898 30.22 / 0.8802 30.83 / 0.8880 30.12 / 0.8808
Table 2: Performance comparison on the C-STVSR baselines for out-of-distribution scale on GoPro dataset. $\mathcal{L}_{RAFT}$ refers the optical flow supervision. Results are evaluated using PSNR (dB) and SSIM metrics. All frames are interpolated by a scaling factor specified on the table and metrics calculated across all interpolated frames. Red and blue indicate the best and the second best performance, respectively.
Temporal Scale Spatial Scale RIFE [13] EMA-VFI [45] VideoINR [7] MoTIF [6] BF-STVSR $+\mathcal{L}_{RAFT}$ (Ours) BF-STVSR (Ours)
LIIF [5] LTE [16] LIIF [5] LTE [16]
×$\times 8$$\times 4$ 29.14 / 0.8524 29.14 / 0.8524 29.68 / 0.8671 29.68 / 0.8667 29.41 / 0.8669 30.04 / 0.8773 30.20 / 0.8799 30.22 / 0.8802
×$\times 4$ 30.16 / 0.8738 30.16 / 0.8737 30.64 / 0.8850 30.64 / 0.8848 30.78 / 0.8954 31.56 / 0.9064 31.68 / 0.9082 31.70 / 0.9083
×$\times 6$$\times 6$ 27.87 / 0.8038 27.86 / 0.8031 28.17 / 0.8126 28.17 / 0.8117 25.56 / 0.7671 29.36 / 0.8505 29.44 / 0.8516 29.45 / 0.8520
×$\times 12$ 24.74 / 0.7019 24.70 / 0.6994 24.85 / 0.7052 24.82 / 0.7028 24.02 / 0.6900 25.81 / 0.7330 25.78 / 0.7284 25.80 / 0.7295
×$\times 4$ 27.43 / 0.8102 27.42 / 0.8100 27.90 / 0.8263 27.90 / 0.8260 27.32 / 0.8141 27.77 / 0.8230 28.06 / 0.8287 28.07 / 0.8287
×$\times 12$$\times 6$ 26.19 / 0.7640 26.19 / 0.7636 26.49 / 0.7748 26.49 / 0.7743 24.68 / 0.7358 26.78 / 0.7908 27.06 / 0.7961 27.07 / 0.7963
×$\times 12$ 24.03 / 0.6869 24.00 / 0.6853 24.16 / 0.6918 24.15 / 0.6902 23.70 / 0.6830 24.72 / 0.7108 24.87 / 0.7096 24.88 / 0.7104
×$\times 4$ 26.08 / 0.7735 26.08 / 0.7733 26.56 / 0.7904 26.56 / 0.7902 25.81 / 0.7739 25.98 / 0.7758 26.40 / 0.7844 26.39 / 0.7840
×$\times 16$$\times 6$ 25.24 / 0.7394 25.24 / 0.7391 25.54 / 0.7503 25.55 / 0.7499 23.86 / 0.7123 25.34 / 0.7527 25.81 / 0.7621 25.81 / 0.7619
×$\times 12$ 23.57 / 0.6781 23.56 / 0.6769 23.68 / 0.6828 23.69 / 0.6816 22.88 / 0.6659 23.88 / 0.6923 24.22 / 0.6950 24.22 / 0.6955